OpenStudy (anonymous):

find the indicated outputs for: f(x)=3xsquared-3x f(0) f(-1) f(2) I have 0,6,6????

OpenStudy (anonymous):

x(3x - 3) f(0) = 0 f(-1) = -(-3-3) f(-1) = 6 f(2) = 2(3(2) - 3)) f(2) = 2(3) f(2) = 6 yep you're right

OpenStudy (anonymous):

thnk u

OpenStudy (anonymous):

f(x) = 3x (x -1) f(0) = 0 f(-1) = -3 * -2 = 6 f(2) = 6